# Mapping 1,000+ Language Models via the Log-Likelihood Vector

### v1

@@ Appendix I Table of Nearest Neighbor Models
767 SeaLLMs/SeaLLM-7B-v2 Kojima et al. (2023); Zheng et al. (2023); Zhang et al. (2023); Nguyen et al. (2024) mistral 7B 2024-01-29 6294 -548.56 67.57
@@
783 speakleash/Bielik-7B-Instruct-v0.1 Levine et al. (2020); Granziol et al. (2021); Ociepa et al. (2024b); Wang et al. (2024a); Ociepa et al. (2024c) mistral 7B 2024-03-30 3573 -871.06 51.24
@@
787 stabilityai/StableBeluga-13B Touvron et al. (2023b); Mukherjee et al. (2023) llama 13B 2023-07-27 6154 -540.85 57.05
788 stabilityai/StableBeluga-7B Touvron et al. (2023b); Mukherjee et al. (2023) llama 6B 2023-07-27 6691 -565.19 53.56
@@
790 stabilityai/stablelm-base-alpha-7b-v2 Shazeer (2020); Rajbhandari et al. (2020); Gao et al. (2020); Li et al. (2023a); Su et al. (2023a); Tow (2023) stablelm_alpha 6B 2023-08-04 2209 -505.45 46.18
@@
792 stabilityai/stablelm-tuned-alpha-3b Taori et al. (2023); Anand et al. (2023); Chiang et al. (2023) gpt_neox 3B 2023-04-19 2145 -736.44 32.14
793 stabilityai/stablelm-tuned-alpha-7b Taori et al. (2023); Anand et al. (2023); Chiang et al. (2023) gpt_neox 7B 2023-04-19 3765 -694.37 34.04
@@
798 stabilityai/stablelm-3b-4e1t Ba et al. (2016); Zhang and Sennrich (2019); Rajbhandari et al. (2020); Gao et al. (2020); Black et al. (2022); Li et al. (2023a); Su et al. (2023a); Tow et al. (2023); Touvron et al. (2023b) stablelm 2B 2023-09-29 11112 -510.56 46.58
799 stabilityai/stable-code-3b Rajbhandari et al. (2020); Black et al. (2022); Li et al. (2023a); Su et al. (2023a); Touvron et al. (2023b); Pinnaparaju et al. (2024); Azerbayev et al. (2024); Yu et al. (2024b) stablelm 2B 2024-01-09 5670 -616.25 41.53
@@
842 TheBloke/stable-vicuna-13B-HF von Werra et al. (2023); Anand et al. (2023); Chiang et al. (2023); Taori et al. (2023); Touvron et al. (2023a) llama-1 13B 2023-04-28 1337 -590.21 51.64
843 TheBloke/tulu-13B-fp16 Köpf et al. (2023); Conover et al. (2023); Wang et al. (2023c); Peng et al. (2023a); Longpre et al. (2023); Chaudhary (2023); Touvron et al. (2023a) llama-1 13B 2023-06-10 1231 -583.06 53.58
844 TheBloke/tulu-7B-fp16 Köpf et al. (2023); Conover et al. (2023); Wang et al. (2023c); Peng et al. (2023a); Longpre et al. (2023); Chaudhary (2023); Touvron et al. (2023a) llama-1 7B 2023-06-10 4079 -616.51 50.24
@@
862 tiiuae/falcon-7b-instruct Shazeer (2019); Brown et al. (2020); Dao et al. (2022); Su et al. (2023a); Penedo et al. (2023); Almazrouei et al. (2023) falcon 7B 2023-04-25 179952 -621.07 43.16
863 tiiuae/falcon-7b Shazeer (2019); Brown et al. (2020); Gao et al. (2020); Dao et al. (2022); Su et al. (2023a); Penedo et al. (2023); Almazrouei et al. (2023) falcon 7B 2023-04-24 104010 -549.44 44.17
@@
993 yanolja/EEVE-Korean-Instruct-10.8B-v1.0 Lian et al. (2023g); Mukherjee et al. (2023); Kim et al. (2024c); Cui et al. (2024) llama 10B 2024-02-22 13241 -555.03 66.48

### v2

@@ Appendix I Table of Nearest Neighbor Models
783 speakleash/Bielik-7B-Instruct-v0.1 Levine et al. (2020); Granziol et al. (2021); Ociepa et al. (2024c); Wang et al. (2024a); Ociepa et al. (2024b) mistral 7B 2024-03-30 3573 -871.06 51.24
@@
787 stabilityai/StableBeluga-13B Mukherjee et al. (2023); Touvron et al. (2023b) llama 13B 2023-07-27 6154 -540.85 57.05
788 stabilityai/StableBeluga-7B Mukherjee et al. (2023); Touvron et al. (2023b) llama 6B 2023-07-27 6691 -565.19 53.56
@@
791 stabilityai/stablelm-base-alpha-7b-v2 Gao et al. (2020); Shazeer (2020); Rajbhandari et al. (2020); Su et al. (2023a); Li et al. (2023a); Tow (2023) stablelm_alpha 6B 2023-08-04 2209 -505.45 46.18
792 stabilityai/stablelm-tuned-alpha-3b Taori et al. (2023); Chiang et al. (2023); Anand et al. (2023) gpt_neox 3B 2023-04-19 2145 -736.44 32.14
793 stabilityai/stablelm-tuned-alpha-7b Taori et al. (2023); Chiang et al. (2023); Anand et al. (2023) gpt_neox 7B 2023-04-19 3765 -694.37 34.04
@@
798 stabilityai/stablelm-3b-4e1t Ba et al. (2016); Zhang and Sennrich (2019); Gao et al. (2020); Rajbhandari et al. (2020); Black et al. (2022); Su et al. (2023a); Tow et al. (2023); Li et al. (2023a); Touvron et al. (2023b) stablelm 2B 2023-09-29 11112 -510.56 46.58
799 stabilityai/stable-code-3b Rajbhandari et al. (2020); Black et al. (2022); Su et al. (2023a); Li et al. (2023a); Touvron et al. (2023b); Yu et al. (2024b); Azerbayev et al. (2024); Pinnaparaju et al. (2024) stablelm 2B 2024-01-09 5670 -616.25 41.53
@@
842 TheBloke/stable-vicuna-13B-HF von Werra et al. (2023); Touvron et al. (2023a); Anand et al. (2023); Taori et al. (2023); Chiang et al. (2023) llama-1 13B 2023-04-28 1337 -590.21 51.64
843 TheBloke/tulu-13B-fp16 Chaudhary (2023); Conover et al. (2023); Touvron et al. (2023a); Wang et al. (2023c); Köpf et al. (2023); Longpre et al. (2023); Peng et al. (2023a) llama-1 13B 2023-06-10 1231 -583.06 53.58
844 TheBloke/tulu-7B-fp16 Chaudhary (2023); Conover et al. (2023); Touvron et al. (2023a); Wang et al. (2023c); Köpf et al. (2023); Longpre et al. (2023); Peng et al. (2023a) llama-1 7B 2023-06-10 4079 -616.51 50.24
@@
862 tiiuae/falcon-7b Shazeer (2019); Brown et al. (2020); Gao et al. (2020); Dao et al. (2022); Su et al. (2023a); Almazrouei et al. (2023); Penedo et al. (2023) falcon 7B 2023-04-24 104010 -549.44 44.17
863 tiiuae/falcon-7b-instruct Shazeer (2019); Brown et al. (2020); Dao et al. (2022); Su et al. (2023a); Almazrouei et al. (2023); Penedo et al. (2023) falcon 7B 2023-04-25 179952 -621.07 43.16
@@
993 yanolja/EEVE-Korean-Instruct-10.8B-v1.0 Mukherjee et al. (2023); Lian et al. (2023g); Kim et al. (2024c); Cui et al. (2024) llama 10B 2024-02-22 13241 -555.03 66.48